Find the value of the following
54279 × 92 + 8 × 54279 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given
54279 × 92 + 8 × 54279
= 54279 × 92 + 54279 × 8
= 54279 × (92 + 8)
= 54279 × 100
= 5427900

Thanks